OpenStudy (anonymous):

FAN AND MEDAL!!!! The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3π (x - 7)2 (x - 3) What does the factor (x - 7)2 represent? The square of the radius of the circular base of the tank The square of the height of the tank The area of the circular base of the tank The area of the curved sides of the tank

OpenStudy (kl0723):

so volume for a tank with a cone shape is: \[\pi r^2 \frac{ h }{ 3 }\]

OpenStudy (kl0723):

now the radius is 4 feet less than the height, if the height is "x-3" would u agree that the radius would be x-3-4?

OpenStudy (jjuden):

The area of the circular base of the tank

OpenStudy (anonymous):

I'm kind of having trouble would you mind helping explain?

OpenStudy (kl0723):

ok, it says that the height is x-3 so if the radius is 4 feet less than the height then you need to take away 4 units from the height to get the radius, so we solve x-3-4= (x-7)

OpenStudy (kl0723):

now, do you see the volume formula for the cone? you have that the radius needs to be squared so if you do that you have: \[(x-7)^2\]

OpenStudy (kl0723):

so if you take a look back in your question, that expression is nothing more than the square of the radius for your cone.

OpenStudy (anonymous):

thank you so much you're even better then my teacher at helping explain problems

OpenStudy (kl0723):

np :)